The order is for nitroprusside 0.3 mcg / kg / min. Titrate upward by 0.1 mcg / kg / min every 2 minutes to a maximum of 2 mcg / kg / min. The patient weighs 165 lb. The pharmacy sends an IV of 50 mg nitroprusside in 500 mL D5W. How many mL / hr can the nurse set on the IV pump to deliver the ordered dosage range of the nitroprusside?
Answer #1

Formula to find ml/hr: [Ordered dose(mcg/kg/min) * patient's weight(kg) * 60 (min/hr)]

___________________________________________________

Medication concentration(mcg/ml)

ordered dose= 0.3 mcg/kg/min

patient weight=165 lbs converting to kg (1 pound =0.45kg) 165*0.45=75 kg

medication concentration=50 mg in 500 ml ;To find per ml concentration=50/500=0.1mg/ml;converting mg to mcg=(1 mg =1000mcg) hence 0.1*1000=100 mcg/ml

By applying the above formula=(0.3 * 75 * 60) / 100=13.5 ml / hr

So the nurse can set 13.5 ml/hr in the infusion pump